OpenStudy (anonymous):

solve for variable x in the following equation pls someone: 2/x-4 + 1/x+4 = 5x/x^2-16 the solution x is

OpenStudy (anonymous):

can someone help me plse: solve for variable x in the following equation pls someone: 2/x-4 + 1/x+4 = 5x/x^2-16 the solution x is

OpenStudy (anonymous):

x is 2. you want the steps?

OpenStudy (anonymous):

yes pls that would be great satellite

OpenStudy (anonymous):

Can you help with this one satellite pls x+1/x+6 = x-1/x=3

OpenStudy (anonymous):

idea is to add on the left hand side. you get \[\frac{2}{x-4}+\frac{1}{x+4}=\frac{2(x+4)+1(x-4)}{x^2-16}\] \[=\frac{3x+4}{x^2-16}\] and the right hand side is \[\frac{5x}{x^2-16}\] so set \[3x+4=5x\] and solve for x via \[3x+4=5x\] \[2x=4\] \[x=2\]

OpenStudy (anonymous):

correction satellite the equation is: x+1/x+6 = x-1/x+3

OpenStudy (anonymous):

that is what i thought. maybe this time we "cross multiply" to get \[(x+1)(x+3)=(x+6)(x-1)\] \[x^2+4x+3=x^2+5x-6\] \[4x+3=5x-6\] \[3+6=5x-4x\] \[9=x\]

OpenStudy (anonymous):

thanks satellite you are very smart have a good night